# Improving Lesion Segmentation in Medical Images by Global and Regional Feature Compensation

### v1

@@ 3.4 Deployment and Objective Function
The GCU and RCU modules can be seamlessly integrated into any U-shaped CNN architecture that utilizes skip connections. Both modules offer flexibility and can be incorporated at various skip concatenation points or decoder layers, adapting to the specific needs of the backbone architecture and dataset characteristics. For the skip connections in the UNet++ backbone, the residuals from the first skip feature are added to subsequent skip connections, as illustrated in Fig. 2. The RCU generates a weighted feature map, which replaces the original upsampled feature map at the corresponding decoder layer.
The framework operates sequentially, with the MAE model first generating MRM from the input images. These MRMs are then used as additional information for the segmentation network, alongside the raw input images. The images are processed slice-wise, and the final output is a single-channel probability map of the same lateral resolution as the input images. A sigmoid activation function is applied to the output, producing per-pixel probabilities for the lesion and the background classification. The patch size of the MRM and the feature maps used in cross-attention are treated as tunable hyper-parameters, allowing optimization for different datasets. The MAE is trained using the Mean Squared Error (MSE) loss as default. For the segmentation network, a mixed loss function combining dice loss [41] and focal loss [42] is utilized for the ATLAS dataset, while focal loss is applied for the orCaScore dataset. This approach balances segmentation precision and recall, enhances sensitivity to challenging lesions, and addresses class imbalance. The mixed loss and focal loss are defined as follows:
@@
This integration of the GCU and RCU into a U-shaped CNN backbone enhances the network’s ability to capture fine-grained details, improving sensitivity and segmentation performance for general segmentation tasks, especially for subtle and small lesions. The flexible design can easily be adaptable to other encoder-decoder backbone architectures, making the approach widely applicable.
@@ 4.1 Datasets
We used two publicly available medical image segmentation datasets as examples to evaluate lesions with low-contrast and small-size characteristics: the Anatomical Tracings of Lesions After Stroke (ATLAS) v2.0 [43] dataset and the Coronary Artery Calcium Scoring (orCaScore) [44] dataset.
@@
To evaluate segmentation performance on small lesions, we utilized the orCaScore dataset from the MICCAI 2014 Challenge on Automatic Coronary Calcium Scoring includes CT scans from 72 patients across four European academic hospitals, acquired using scanners from different vendors. Each patient has a non-contrast enhanced, ECG-triggered cardiac calcium scoring CT (CSCT) scan. Coronary artery calcifications (CAC) were manually annotated by an experienced radiologist and a physician. Annotations were provided for calcifications with in-plane resolutions of 0.4–0.5 mm and slice thicknesses/spacings of 2.5–3 mm. The dataset is divided into 32 patients for training and 40 patients for testing (original set). However, GT masks are only available for the training set and not for the testing set. Therefore, we split the 32 patients in the training set into 26 for training, 5 for validation, and 1 for testing. In this study, after model building, the evaluation was performed using the original set of 40 patients for testing, by submitting segmentation results to the official dataset evaluation protocol  [45].
@@ 4.2 Evaluation Setup
We used UNet++ as our segmentation model baseline due to its enhanced performance compared to U-Net in the two segmentation tasks evaluated in this study. To evaluate the performance of the GRCSF, we compared it against several state-of-the-art CNN and transformer-based methods commonly used in medical image segmentation: U-Net and UNet++, which utilize skip connection to enhance multi-scale features; Deeplabv3, which introduces various receptive fields to obtain better spatial information; TransUNet and SwinUNet, which take advantage of self-attention mechanism for better global context; and SPiN which is specifically developed for stroke lesion segmentation. All methods were trained and tested on the same datasets for a fair comparison. The comparison results for the ATLAS and orCaScore datasets are presented in Table 1.
@@
To investigate the contributions of different components in GRCSF, we conducted ablation studies using the ATLAS dataset and UNet++ as the backbone architecture. These experiments aimed to assess the performance improvements introduced by the proposed modules of GCU, RCU and Importance Scoring mechanism. The GCU was applied at each skip connection layer in the backbone, although it can be adjusted depending on the task. Similarly, the RCU was applied to the last three upsampling layers in this study, but it can be selectively disconnected from specific layers as required. The results are summarized in Table 2.
@@
For the ATLAS dataset, the four-subset design was implemented to evaluate the generalizability of the models across different medical centers and scanners while using relatively small amounts of training data compared to perform cross-validation on the entire dataset. The primary performance metric for the ATLAS dataset segmentation was the Dice Similarity Coefficient (DSC), which quantifies the similarity between the predicted and GT regions. Additionally, the Intersection over Union (IoU) was used to measure the overlap between the predicted and GT regions, while the F1 score highlighted the balance between precision and recall. Precision and recall scores were also reported. All metrics were computed on a pixel-wise basis per patient to ensure precision in the evaluation.
@@
The orCaScore evaluation framework is a publicly accessible, web-based protocol for algorithm evaluation, originally introduced at the MICCAI 2014 workshop in Boston, USA. This framework evaluates algorithms based on the number and volume of identified calcifications on a per-patient basis. For this study, the test results were post-processed according to specific guidelines: coronary calcifications were defined as connected voxel groups (using 3D 6-connectivity) with intensities exceeding 130 Hounsfield Units (HU) [46].
@@ 4.3 Implementation Details
For the ATLAS dataset, the 655 cases were divided into four subsets, which is similar to the data split in [47], with each subset containing patients from distinct cohorts. Within each subset, patients were further split into training and testing sets, with 80% allocated for training and 20% for testing. In particular, test patients were drawn from cohorts different from those in the training set. The pre-trained MAE model, enhanced with an additional GAN loss for more realistic generation (ViT-Large architecture, training mask ratios of 50% and 75%), was trained using all non-lesion training slices from all subsets. The training process was conducted over 500 epochs with a batch size of 64. The Adam optimizer [48] was utilized to minimize MSE loss, with an initial learning rate of 0.001. The learning rate decayed by a factor of 0.9 every 50 epochs to ensure stable convergence. The segmentation model was trained with an input size of 233x197 using the early stopping strategy, which stops training if the validation loss did not decrease for 15 consecutive epochs. A batch size of 21 was used, and the Adam optimizer was employed with an initial learning rate of 0.0001. Optimization incorporated a cosine warmup schedule [49] during the first 5 epochs, where the learning rate increased 10 times. The first and second moment estimates for the optimizer were set to 0.9 and 0.999, respectively. The weights of the convolutional filter were initialized using the He method [50], and the biases were initialized to zero. For the cross-attention operation in the RCU module, the feature patch sizes at the last three decoder layers were configured as 16×16, 16×16 and 32×32, respectively.
For the orCaScore dataset, following the data split described in the method section, the same pre-trained MAE models were trained using all non-lesion training slices over 200 epochs with a batch size of 8. The Adam optimizer was employed to minimize the MSE loss, starting with an initial learning rate of 0.001, which decayed by a factor of 0.9 every 50 epochs. The segmentation network training followed the same configuration as for the ATLAS dataset, with the following adjustments: an input size of 512×512, a smaller batch size of 2 due to computational resource constraints and a reduced initial learning rate of 0.0001. Additionally, the learning rate decayed by a factor of 0.5 if the validation loss did not decrease over 5 consecutive epochs. The RCU was presented in the last three decoder layers, the feature patch sizes were configured as 16×16, 32×32 and 32×32, respectively.
@@ 4.4.1 State of the Art comparisons
For the ATLAS dataset, our GRCSF achieved a DSC of 0.422, outperforming the previous state-of-the-art method DeepLabv3, which ranked second, by 2.7% (Table 1). Additionally, our method demonstrated dominant performance across other metrics, achieving an IoU of 0.319 and an F1 Score of 0.473, representing improvements of 1.5% and 0.8% over the second-highest metrics achieved by TransUnet for IoU and DeepLabv3 for F1 Score, respectively.
Methods ATLAS orCaScore
Dice IOU F1 Precision Recall F1 volume Sens vol PPV vol Sen lesion PPV lesion
UNet 0.372 0.276 0.430 0.418 0.444 0.924 0.971 0.881 0.871 0.841
UNet++ 0.381 0.286 0.440 0.409 0.476 0.937 0.954 0.920 0.801 0.880
DeepLabv3 0.395 0.295 0.465 0.413 0.531 0.936 0.929 0.944 0.516 0.925
SwinUnet 0.281 0.204 0.338 0.271 0.450 0.370 0.251 0.701 0.275 0.584
TransUnet 0.394 0.304 0.440 0.455 0.427 0.770 0.673 0.902 0.149 0.840
SPiN 0.376 0.288 0.427 0.404 0.453 0.913 0.896 0.930 0.253 0.892
GRCSF (Ours) 0.422 0.319 0.473 0.451 0.497 0.946 0.962 0.930 0.827 0.899
Table 1: Comparison of seven segmentation methods on the test sets of ATLAS and orCaScore datasets. The best results are highlighted in bold, and the second-best results are underlined. The metrics for ATLAS represent the average across four test subsets, while the metrics for orCaScore are derived from the online evaluation framework.
Methods Dice IOU F1 Precision Recall
UNet++ 0.381 0.286 0.440 0.409 0.476
GRCSF w/o RCU 0.394 0.294 0.453 0.426 0.485
GRCSF w/o Importance Score in RCU 0.408 0.306 0.457 0.439 0.478
GRCSF 0.422 0.319 0.473 0.451 0.497
@@
Selected example segmentation results from ATLAS and orCaScore images are shown in Fig. 6 and Fig. 7. The ATLAS results in Fig. 6 demonstrate the effectiveness of GCU and RCU, as most methods tended to under-segment for challenging cases, whereas our method successfully captured the structures that other methods missed. The orCaScore results include both post-processed and non-post-processed visualizations of the same prediction, providing additional insights into model performance. In Fig. 7a, the post-processed results of all methods demonstrate accurate lesion boundaries in certain methods; however, this improvement is partly due to the removal of calcifications with image pixel values below 130HU during post-processing. This artificially improves the appearance of over-segmentation methods like U-Net and DeepLabv3, masking their inherent limitations. In contrast, the non-post-processed results in Fig. 7b reveal GRCSF’s true advantage of accurately segmenting small calcifications (as highlighted by the green arrows in Fig. 7) without significant over-segmentation. This highlights our model’s ability to capture challenging structures while maintaining recall without relying on post-processing.
@@
The impact of integrating MRM on segmentation performance is illustrated in Fig. 8. The segmentation of the challenging lesion improved consistently when MRM was incorporated, particularly when using UNet++ as the backbone. This configuration showed the best performance across both datasets.
@@
Figure 7: Visual comparison of our proposed method GRCSF, against previous state-of-the-art methods in medical image segmentation using the orCaScore dataset. (a) Visualization of predictions after post-processing, leveraging domain knowledge that calcifications are identified as regions where the predicted image pixel values exceed 130 HU. (b) Visualization of raw predictions without post-processing. Please zoom in for a clearer view.
Figure 8: Example outputs from UNet++ and our proposed method GRCSF for the ATLAS dataset. The figure includes GT segmentations and the corresponding MRM of the input images for comparison. Please zoom in for a clearer view.

### v2

@@ 3.4 Deployment and Objective Function
The GCU and RCU modules can be integrated into any U-shaped CNN architecture that utilizes skip connections. Both modules offer flexibility and can be incorporated at various skip concatenation points or decoder layers, adapting to the specific needs of the backbone architecture and dataset characteristics. For the skip connections in the UNet++ backbone, the residuals from the first skip feature are added to subsequent skip connections (Fig. 2). The RCU generates a weighted feature map, which replaces the original upsampled feature map at the corresponding decoder layer.
The framework operates sequentially, with the MAE model first generating MRM from the input images. These MRMs are then used as additional information for the segmentation network, alongside the raw input images. The images are processed slice-wise, and the network outputs a single-channel probability map matching the input resolution after a sigmoid activation. The patch size of the MRM and the feature maps used in cross-attention are treated as tunable hyper-parameters. The MAE is trained using the mean squared error (MSE) loss by default. For the segmentation network, a mixed loss function combining dice loss and focal loss is utilized for the ATLAS and the MSD Lung Tumor datasets, while focal loss is applied for the orCaScore dataset. This approach balances segmentation precision and recall, enhances sensitivity to challenging lesions, and addresses class imbalance. The mixed loss and focal loss are defined as follows:
@@
This integration of the GCU and RCU into a U-shaped CNN backbone enhances the network’s ability to capture fine-grained details, improving sensitivity to subtle and small lesions. The flexible design can easily be adaptable to other encoder-decoder backbone architectures, making the approach widely applicable.
@@ 4.1 Datasets
We used three publicly available medical image segmentation datasets as examples to evaluate lesions with low-contrast and small-size characteristics: the Anatomical Tracings of Lesions After Stroke (ATLAS) v2.0 [38] dataset, the Medical Segmentation Decathlon (MSD) Lung Tumor Segmentation dataset [39] and the Coronary Artery Calcium Scoring (orCaScore) [40] dataset.
@@
To evaluate segmentation performance on small lesions, we utilized the MSD Lung Tumor dataset and the orCaScore dataset. The MSD Lung Tumors dataset is acquired from The Cancer Imaging Archive (TCIA), which focuses on the segmentation of NSCLC lesions from CT scans. The dataset consists of thin-section CT scans from 96 patients with NSCLC, officially divided into 64 cases for training and 32 cases for testing. In our study, we used the 64 cases with publicly available CT images and corresponding ground truth masks. The dataset was further divided into 54 patients for training and 10 patients for testing. The training set was then split into 43 patients for training and 11 patients for validation. All scans are non-contrast-enhanced chest CTs, acquired using various scanner settings. The in-plane resolution of the CT scans ranges from 0.60 mm to 0.98 mm, with slice thicknesses varying between 0.63 mm and 2.5 mm. We also included the orCaScore dataset from the MICCAI 2014 Challenge on Automatic Coronary Calcium Scoring contains CT scans from 72 patients across four European academic hospitals, acquired using scanners from different vendors. Each patient has a non-contrast enhanced, ECG-triggered cardiac calcium scoring CT (CSCT) scan. Coronary artery calcifications (CAC) were manually annotated by an experienced radiologist and a physician. Annotations were provided for calcifications with in-plane resolutions of 0.4–0.5 mm and slice thicknesses/spacings of 2.5–3 mm. The dataset is divided into 32 patients for training and 40 patients for testing (original set). However, GT masks are only available for the training set and not for the testing set. Therefore, we split the 32 patients in the training set into 26 for training, 5 for validation, and 1 for testing. In this study, after model building, the evaluation was performed using the original set of 40 patients for testing, by submitting segmentation results to the official dataset evaluation protocol  [41].
@@ 4.2 Evaluation Setup
We used UNet++ as our segmentation model baseline due to its enhanced performance compared to U-Net in the three segmentation tasks evaluated in this study. To evaluate the performance of the GRCSF, we compared it against several state-of-the-art CNN and transformer-based methods commonly used in medical image segmentation: U-Net and UNet++, which utilize skip connection to enhance multi-scale features; Deeplabv3, which introduces various receptive fields to obtain better spatial information; TransUNet, SwinUNet, TransFuse and UTNet, which take advantage of self-attention mechanism for better global context; and SPiN which is specifically developed for stroke lesion segmentation. We also compared our method with the foundation model SAM, which benefits from pre-training on a large-scale dataset. For a fair comparison, all methods were trained and tested on the same datasets, except for SAM, which was evaluated in a zero-shot setting using bounding box prompts derived from UNet++ predictions. The comparison results on the ATLAS and MSD Lung Tumor datasets are shown in Table 1, while the results for the orCaScore dataset are presented in Table 2.
@@
To investigate the contributions of different components in GRCSF, we conducted ablation studies using the ATLAS dataset and UNet++ as the backbone. These experiments aimed to assess the performance improvements introduced by the proposed modules of GCU, RCU and Importance Scoring mechanism. The GCU was applied at each skip connection layer in the backbone, although it can be adjusted depending on the task. Similarly, the RCU was applied to the last three upsampling layers in this study, but it can be selectively disconnected from specific layers as required. The results are summarized in Table 3. We also validated key hyperparameter choices, including the number of MAE iterations for generating residual maps, the associated inference time, the selection of MAE mask ratios, and the patch size for cross-attention in the RCU module. These experiments were conducted on one training and test subset of the ATLAS dataset.
In addition, we performed a comprehensive analysis of alternative design choices within each GRCSF module. This includes evaluating different backbone architectures, such as the lightweight MobileNet [42]. For the GCU, we explored variants using spatial attention modules (SAM) [43] instead of SE blocks, and compared placement strategies involving pyramidal attention modules (PAM) [44] and attention-gated (AG) [45] skip connections. We also assessed different strategies for residual map construction, including Grad-CAM maps generated from UNet predictions, an alternative reconstruction-based method SimMIM [20], and different methods for computing the difference between the original and reconstructed images, including AbsDiff, MSE, and structural similarity index (SSIM). To demonstrate the advantage of reconstruction over simple pre-training, we additionally evaluated a model that uses a MAE-pretrained encoder to replace the original UNet encoder. The results are presented in Table 7.
@@
For the ATLAS dataset, the four-subset design was implemented to evaluate the generalizability of the models across different medical centers and scanners while using relatively small amounts of training data compared to perform cross-validation on the entire dataset. The primary performance metric for both the ATLAS and MSD Lung Tumor dataset was the Dice Similarity Coefficient (DSC), which quantifies the similarity between the predicted and GT regions. Intersection over Union (IoU) was also reported to measure the overlap between the predicted and GT regions. Additional metrics include precision, recall, false positive rate (FPR), and volume over-segmentation error (VOSE). FPR and VOSE quantify the degree of over-segmentation relative to the background and ground-truth lesion volume, respectively. All metrics were computed on a per-patient, pixel-wise basis to ensure precise evaluation.
@@
The orCaScore evaluation framework is a publicly accessible, web-based protocol for algorithm evaluation, originally introduced at the MICCAI 2014 workshop in Boston, USA. This framework evaluates algorithms based on the number and volume of identified calcifications on a per-patient basis. For this study, the test results were post-processed according to specific guidelines: coronary calcifications were defined as connected voxel groups (using 3D 6-connectivity) with intensities exceeding 130 Hounsfield Units (HU). To enable a more comprehensive comparison, we also report pre- and post-processed metrics using a held-out test patient separated from the training set.
@@ 4.3 Implementation Details
For the ATLAS dataset, the 655 cases were divided into four subsets, which is similar to the data split in [46], with each subset containing patients from distinct cohorts. Within each subset, patients were further split into training and testing sets, with 80% allocated for training and 20% for testing. In particular, test patients were drawn from cohorts different from those in the training set. The pre-trained MAE model, enhanced with an additional GAN loss for more realistic generation (ViT-Large architecture, training mask ratios of 50% and 75%), was trained using all non-lesion training slices from all subsets. The training process was conducted over 200 epochs with a batch size of 32. The Adam optimizer was utilized to minimize MSE loss, with an initial learning rate of 0.001. The learning rate decayed by a factor of 0.9 every 50 epochs to ensure stable convergence. The segmentation model was trained with an input size of 224x224 using the early stopping strategy, which stops training if the validation loss did not decrease for 10 consecutive epochs. A batch size of 21 was used, and the Adam optimizer was employed with an initial learning rate of 0.0001. Optimization incorporated a cosine warmup schedule during the first five epochs, where the learning rate increased 10 times. The first and second moment estimates for the optimizer were set to 0.9 and 0.999, respectively. The weights of the convolutional filter were initialized using the He method [47], and the biases were initialized to zero. For the cross-attention operation in the RCU module, the feature patch sizes at the last three decoder layers were configured as 8×8, 8×8 and 16×16, respectively. The MSD Lung Tumor dataset used the data split strategy described in the Datasets section and was trained under the same settings as the ATLAS dataset.
For the orCaScore dataset, following the data split described in the Datasets section, the same pre-trained MAE models were trained using all non-lesion training slices over 200 epochs with a batch size of 32. The Adam optimizer was employed to minimize the MSE loss, starting with an initial learning rate of 0.001, which decayed by a factor of 0.9 every 50 epochs. The segmentation network training followed the same configuration as for the ATLAS dataset, with the following adjustments: an input size of 512×512, a smaller batch size of 2 due to computational resource constraints and a reduced initial learning rate of 0.0001. Additionally, the learning rate decayed by a factor of 0.5 if the validation loss did not decrease over 5 consecutive epochs. The RCU was presented in the last three decoder layers, the feature patch sizes were configured as 16×16, 16×16 and 32×32, respectively.
@@
Methods ATLAS MSD Lung Tumor Model Complexity
Dice IoU Precision Recall VOE FPR Inference Time Dice IoU Precision Recall VOE FPR Inference Time Model Parameters GFLOPs Peak Memory
UNet 0.372 0.276 0.444 0.418 0.818 0.047% 2.42s 0.670 0.521 0.724 0.702 0.528 0.005% 25.62s 31.04M 41.85 210.7MB
UNet++ 0.381 0.286 0.476 0.409 0.914 0.054% 3.48s 0.691 0.538 0.643 0.770 0.471 0.010% 36.63s 36.63M 106.1 327.8MB
DeepLabv3 0.395 0.295 0.531 0.413 0.485 0.038% 3.36s 0.633 0.473 0.660 0.672 0.525 0.006% 37.28s 39.63M 31.41 301.1MB
SwinUnet 0.281 0.204 0.450 0.271 0.703 0.038% 3.86s 0.248 0.148 0.475 0.198 0.387 0.005% 36.84s 41.34M 8.69 212.5MB
TransUnet 0.394 0.304 0.427 0.450 0.871 0.069% 4.31s 0.680 0.533 0.657 0.776 0.689 0.006% 37.23s 93.23M 24.67 803.7MB
SPiN 0.376 0.288 0.453 0.404 1.100 0.061% 3.69s 0.683 0.534 0.621 0.832 0.766 0.008% 54.02s 5.17M 6.23 171.9MB
TransFuse 0.366 0.274 0.407 0.476 1.448 0.077% 6.21s 0.586 0.425 0.478 0.800 1.071 0.010% 65.23s 143.39M 63.40 737.6MB
UTNet 0.344 0.259 0.466 0.430 0.465 0.042% 4.10s 0.645 0.500 0.695 0.659 0.296 0.005% 36.18s 80.77M 81.83 515.1MB
SAM 0.360 0.261 0.438 0.399 1.089 0.010% 65.33s 0.722 0.583 0.745 0.728 0.307 0.007% 79.50s 631.58M 10934.57 5746MB
GRCSF (Ours) 0.422 0.319 0.497 0.451 0.942 0.050% 25.93s * 0.730 0.583 0.709 0.780 0.370 0.005% 61.59s * 42.85M 123.10 690.0MB
Table 1: Comparison of ten segmentation methods on the ATLAS and MSD Lung Tumor test sets. The best scores are shown in bold and the second-best are underlined. ATLAS metrics are first averaged across patients within each of the four test subsets and then averaged across the subsets, whereas MSD Lung Tumor metrics are averaged over all test patients. Inference time is reported as the mean per patient. * indicates that the reported time is the end-to-end runtime of the entire framework, including the time required for MAE residual map generation and segmentation. “Model Complexity” lists the number of parameters, Giga floating-point operations per second (GFLOPs), and peak GPU memory consumption.
orCaScore
Methods Test Patients (Post-Processed) Validation Patient (Post-Processed) Validation Patient (without Post-Processing)
F1 vol Sens vol PPV vol Sen lesion PPV lesion F1 vol Sens vol PPV vol VOE F1 vol Sens vol PPV vol VOE
UNet 0.924 0.971 0.881 0.871 0.841 0.889 0.808 0.989 0.009 0.743 0.808 0.687 0.368
UNet++ 0.937 0.954 0.920 0.801 0.880 0.921 0.859 0.992 0.007 0.843 0.859 0.828 0.178
DeepLabv3 0.936 0.929 0.944 0.516 0.925 0.971 0.951 0.992 0.008 0.636 0.951 0.478 1.041
SwinUnet 0.370 0.251 0.701 0.275 0.584 0.467 0.323 0.843 0.060 0.400 0.323 0.525 0.292
TransUnet 0.770 0.673 0.902 0.149 0.840 0.806 0.730 0.900 0.081 0.804 0.730 0.895 0.085
SPiN 0.913 0.896 0.930 0.253 0.892 0.851 0.746 0.990 0.009 0.842 0.746 0.967 0.033
TransFuse 0.749 0.635 0.913 0.269 0.917 0.811 0.682 1.000 0.000 0.513 0.682 0.410 0.981
UTNet 0.859 0.807 0.919 0.234 0.840 0.939 0.886 0.999 0.001 0.825 0.886 0.771 0.263
SAM 0.931 0.971 0.895 0.898 0.842 0.913 0.841 0.999 0.001 0.703 0.841 0.604 0.552
GRCSF (Ours) (Ours) 0.946 0.962 0.930 0.827 0.899 0.885 0.799 0.992 0.006 0.856 0.799 0.921 0.068
Table 2: Comparison of segmentation methods on the orCaScore dataset across three subsets: the 40 online-test patients after post-processing (metrics from the online evaluation framework), a held-out test patient after post-processing, and the same patient without post-processing (metrics computed from its ground-truth mask). The best results are shown in bold and the second-best are underlined.
Methods Dice IoU Precision Recall
UNet++ 0.381 0.286 0.476 0.409
GRCSF w/o RCU 0.394 0.294 0.485 0.426
GRCSF w/o Importance Score in RCU 0.408 0.306 0.478 0.439
GRCSF 0.422 0.319 0.497 0.451
@@ 4.4.1 State-of-the-Art Comparisons
For the ATLAS dataset, our GRCSF achieved a DSC of 0.422, outperforming the previous state-of-the-art method DeepLabv3, which ranked second, by 2.7% (Table 1). Furthermore, our method demonstrated a dominant performance in IoU of 0.319, representing improvements of 1.5% over the second highest metrics obtained by TransUnet. GRCSF achieved the second-highest precision and recall scores, suggesting balanced segmentation performance that neither oversegments nor undersegments compared to the other methods.
For the MSD Lung Tumor dataset, our method achieved a DSC of 0.730, outperforming the foundation model SAM by 0.08% and surpassing UNet++, the third-best performer by 3.9% (Table 1). In terms of IoU, our method reached a score of 0.583, matching that of SAM and exceeding the second-highest UNet++ by 4.5%. Furthermore, GRCSF obtained the lowest FPR, indicating a reduced tendency to over-segment background regions.
In terms of model complexity, the total prediction time per patient for the entire GRCSF framework is 25.93 seconds on the ATLAS dataset and 61.59 seconds on the MSD Lung Tumor dataset. This end-to-end runtime consists of two components: the segmentation network, which requires 4.07 seconds for ATLAS and 34.93 seconds for MSD Lung Tumor; and the residual map generation process, which adds 21.86 seconds (Table 4) and 26.66 seconds, respectively, when using five MAE iterations and two mask ratios. GRCSF has 42.85M parameters, similar to SwinUNet with 41.34M and slightly higher than UNet++ with 36.63M, while remaining lighter than Transformer-based models such as TransUnet with 93.23M and TransFuse with 143.39M. Its GFLOPs reach 123.10, which is higher than those of CNN-based models like UNet (41.85) and DeepLabv3 (31.41), but much lower than that of SAM, which has 10,934.57 GFLOPs. GRCSF’s peak memory usage is 690.0 MB, lower than that of TransUnet (803.7 MB) and TransFuse (737.6 MB). Overall, our method strikes a favorable balance between efficiency and accuracy.
@@
Selected example segmentation results from ATLAS, MSD Lung Tumor and orCaScore are shown in Fig. 6, Fig. 7 and Fig. 8. The visualizations demonstrate the effectiveness of GCU and RCU, as most methods tended to under-segment for challenging cases, whereas our method successfully captured the structures that other methods missed. As shown in Fig. 9, the case-wise performance breakdown by lesion size demonstrates that our method outperforms all others on small and ambiguous lesions ranging from 1 to 1,000 pixels. TransUnet achieves the best performance in the subgroup of lesions larger than 1,000 pixels, however our method ranks the second in that category and when averaged across all the sizes, our method is the best. The orCaScore results include both post-processed and non-post-processed visualizations of the same prediction with error maps, providing additional insights into model performance. In Fig. 8a, the post-processed results of all methods demonstrate accurate lesion boundaries in certain methods; however, this improvement is partly due to the removal of calcifications with image pixel values below 130HU during post-processing. This artificially improves the appearance of over-segmentation methods like U-Net and DeepLabv3, masking their inherent limitations. In contrast, the non-post-processed results in Fig. 8b reveal GRCSF’s true advantage of accurately segmenting small calcifications without significant over-segmentation. This highlights our model’s ability to capture challenging structures while maintaining recall without relying on post-processing.
@@
The impact of integrating MRM on segmentation performance is illustrated in Fig. 10. The segmentation of the challenging lesion improved consistently when MRM was incorporated, particularly when using UNet++ as the backbone. This configuration showed the best performance across three datasets.
@@
MAE Iterations Dice IoU Precision Recall Inference Time
1 0.520 0.400 0.561 0.568 4.33s
2 0.554 0.421 0.557 0.656 8.71s
3 0.534 0.404 0.505 0.638 12.30s
4 0.550 0.417 0.582 0.646 18.24s
5 0.581 0.448 0.595 0.645 21.86s
6 0.573 0.439 0.603 0.636 24.08s
Table 4: Performance of GRCSF with residual maps reconstructed from different numbers of MAE iterations, evaluated on one ATLAS test subset. Inference time is reported as the mean per patient while generating residual maps for two MAE mask ratios. The best results are highlighted in bold.
Experiment 30% 50% 75% 90% Dice IoU Precision Recall
1 0.537 0.413 0.557 0.627
2 0.574 0.439 0.569 0.681
3 0.541 0.408 0.601 0.636
4 0.531 0.403 0.580 0.572
5 0.581 0.448 0.595 0.645
Table 5: Performance of GRCSF using residual maps generated with different MAE mask ratios, evaluated on one ATLAS test subset. The best results are highlighted in bold.
Cross Attention Patch Size Dice IoU Precision Recall Model Parameters
4×4, 8×8, 16×16 0.500 0.373 0.426 0.734 42.75M
8×8, 4×4, 16×16 0.482 0.357 0.408 0.727 42.82M
8×8, 8×8, 8×8 0.480 0.358 0.604 0.492 42.75M
8×8, 16×16, 16×16 0.553 0.415 0.585 0.583 42.95M
8×8, 8×8, 16×16 (Ours) 0.581 0.448 0.595 0.645 42.85M
Table 6: Performance of GRCSF with different patch sizes for cross-attention in the RCU, evaluated on one ATLAS test subset. The best results are highlighted in bold.
@@
Figure 7: Visual comparison of our proposed method GRCSF, against previous state-of-the-art methods in medical image segmentation using the MSD Lung Tumor dataset. Please zoom in for a clearer view.
Figure 8: Visual comparison of our proposed method GRCSF, against previous state-of-the-art methods in medical image segmentation using the orCaScore dataset. (a) Visualization of predictions after post-processing, leveraging domain knowledge that calcifications are identified as regions where the predicted image pixel values exceed 130 HU. (b) Visualization of raw predictions without post-processing. Error maps are included to highlight prediction errors on small lesions more clearly: red indicates false-positive pixels, while green indicates false negatives. Please zoom in for a clearer view.
Figure 9: Case-wise lesion-size performance of all methods on one ATLAS test subset. Please zoom in for a clearer view.
Figure 10: Example outputs from UNet++ and our proposed method GRCSF for the ATLAS and the MSD Lung Tumor datasets. The figure includes GT segmentation and the corresponding MRM of the input images for comparison. Please zoom in for a clearer view.